# KG-Infused RAG: Augmenting Corpus-Based RAG with External Knowledge Graphs

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) improves factual accuracy by grounding responses in external knowledge. However, existing methods typically rely on a single source, either unstructured text or structured knowledge. Moreover, they lack cognitively inspired mechanisms for activating relevant knowledge. To address these issues, we propose KG-Infused RAG, a framework that integrates KGs into RAG systems to implement spreading activation, a cognitive process that enables concept association and inference. KG-Infused RAG retrieves KG facts, expands the query accordingly, and enhances generation by combining corpus passages with structured facts, enabling interpretable, multi-source retrieval grounded in semantic structure. We further improve KG-Infused RAG via preference learning on sampled key stages in the pipeline. Experiments on five QA benchmarks show that KG-Infused RAG consistently outperforms vanilla RAG (by 3.8% to 13.8%). Additionally, when integrated into Self-RAG, KG-Infused RAG brings further performance gains, demonstrating its effectiveness and versatility as a plug-and-play enhancement module for corpus-based RAG methods.  [^1] Our code and data are available at https://github.com/thunlp/KG-Infused-RAG
@@ 1 Introduction
Large language models (LLMs) have shown strong performance in question answering tasks [1, 10, 30, 5], yet they remain susceptible to factual errors and hallucinations [19, 21, 8]. Retrieval-Augmented Generation (RAG) mitigates these issues by grounding generation in external sources such as text corpus or knowledge graphs (KGs) [17, 25, 2, 3, 28]. However, existing RAG methods typically rely on a single retrieval source: corpus-based methods retrieve passages [8, 22, 25, 2], while KG-based methods extract facts from an external graph [3, 28]. This single-source constraint limits their ability to integrate complementary evidence from both unstructured and structured knowledge. Recent hybrid approaches attempt to bridge this gap by constructing ad-hoc KGs from text using LLMs [6, 11, 32], but they often suffer from high computation costs and potential factual inaccuracies due to imperfect entity and relation extraction.
More fundamentally, current RAG methods lack the ability to retrieve knowledge by activating semantically related concepts based on prior knowledge structures. This cognitive process, known as spreading activation [4], originates from cognitive psychology and explains how humans access relevant information by traversing semantic associations, even when it is not explicitly mentioned. In RAG, this mechanism can be implemented through the use of KGs, where a seed concept (e.g., an entity in the question) activates connected nodes (e.g., related entities or facts), guiding the retrieval toward semantically relevant evidence. This enables richer knowledge coverage, improves recall of long-tail or implicit facts, and provides a more interpretable reasoning path compared to relying solely on surface-form matching or parametric memory in LLMs.
To efficiently implement spreading activation, we leverage existing human-curated KGs (e.g., Wikidata), which provide high-quality, structured, and long-tail factual knowledge. Unlike prior work that constructs ad-hoc KGs [6, 11, 32] from text using LLMs—often prone to factual errors and computational costs—those curated KGs offer a stable and cost-effective foundation for semantically guided retrieval. Using KGs as external memory networks enable RAG systems to traverse meaningful concept associations, improving both recall and interpretability of the retrieved evidence.
@@
Based on this insight, we introduce Knowledge Graph Infused Retrieval-Augmented Generation (KG-Infused RAG), a framework that integrates KGs into the RAG pipeline to implement spreading activation. KG-Infused RAG consists of three stages: (1) retrieving relevant KG facts via spreading activation; (2) expanding the query with these facts to improve corpus retrieval; (3) generating answers conditioned on passages enriched with KG facts. This design enables multi-source retrieval that combines KG facts and corpus passages, leveraging KG structure to guide spreading activation. By aligning retrieval with human-like spreading activation, KG-Infused RAG supports more accurate, interpretable, and fact-grounded QA. Moreover, it can be integrated into corpus-based RAG methods (e.g., Self-RAG) to further boost performance.
@@
    We introduce KG-Infused RAG, a framework that incorporates human-curated knowledge graphs (KGs) into RAG to emulate spreading activation for semantically guided retrieval and generation.
@@
    KG-Infused RAG supports multi-source retrieval and can be integrated into corpus-based RAG methods (e.g., Self-RAG) as a plug-and-play module.
@@
    We develop a data sampling method to improve preference learning and model performance.
@@
    Experiments on five QA benchmarks show that KG-Infused RAG consistently improves performance, with absolute gains of 3.8% to 13.8% over vanilla RAG and substantial improvements when combined with Self-RAG.
@@ 2.1 Multi-Turn Retrieval in Retrieval-Augmented Generation
Single-turn retrieval followed by generation often yields insufficient evidence for complex questions, such as multi-hop QA requiring LLMs to perform multi-step retrieval and reasoning [8, 25].
@@
Query rewriting reformulates an original query into a clearer or more effective one, especially when it is ambiguous or complex [18, 20]. Self-Ask [22] and IRCoT [25] enhance multi-hop reasoning by decomposing complex queries into sub-questions for iterative retrieval. In parallel, query expansion enrich the original query with informative content, such as keywords or pseudo-documents, to improve retrieval [26, 14, 16]. HyDE [7] generates hypothetical documents from the query to improve retrieval relevance. In contrast to these methods that rely solely on an LLM’s parametric knowledge, KG-Infused RAG incorporates factual signals from a KG during query expansion. Grounded in relevant facts, KG-Infused RAG enables more targeted and evidence-grounded retrieval.
@@
Recent RAG methods extend single-turn retrieval to multi-turn, collecting richer and more targeted evidence [8]. IRCoT [25] interleaves retrieval with chain-of-thought reasoning steps, while FLARE [15] and Self-RAG [2] dynamically decide when and what to retrieve based on generation confidence or intermediate outputs. These methods enable incremental evidence accumulation during generation. Similarly, KG-Infused RAG gathers additional evidence through multi-turn spreading activation and KG-based query expansion, followed by integration and refinement of retrieved information.
@@ 2.2 Knowledge Graphs Augmented Question Answering
Knowledge graphs (KGs) provide structured representations of entities and their relations, offering signals beyond plain text. This structured knowledge helps reduce hallucinations, improves reasoning interpretability, and enhances retrieval and generation in RAG systems [3, 29, 28, 11]. MindMap [28] and Xu et al. [29] use KGs as retrieval sources instead of text corpus in domains like medical and customer service QA, achieving better retrieval accuracy, answer quality, and transparency. In contrast, GraphRAG [6], LightRAG [11], and KG2RAG [32] generate ad-hoc KGs from the whole corpus or retrieved passages to organize entity relationships and improve answer coherence for complex questions. In contrast, KG-Infused RAG directly leverages existing high-quality, human-curated KGs, avoiding costly construction while providing reliable and structured external knowledge.
@@ 3.1 Task Definition
Let $\mathcal{G}_{q}\subset\mathcal{G}$ denote the retrieved KG subgraph composed of query-relevant triples, and $\mathcal{P}_{q}\subset\mathcal{D}_{c}$ the corresponding retrieved passages. As the KG can be viewed as a semantic network of concept-level associations, the entities in $\mathcal{G}_{q}$ serve as seeds for activation propagation over the KG. Formally, the task can be framed as:
$\text{answer}=\text{LLM}\left(q,\mathcal{G}_{q},\mathcal{P}_{q}\right)$ (1)
@@ 3.2 Framework Overview
Figure 1: An overview of the framework KG-Infused RAG.
We propose KG-Infused RAG, a framework inspired by the cognitive process of spreading activation, where activation spreads from a central concept to semantically associated ones. By leveraging the semantic structure of KGs, KG-Infused RAG expands relevant knowledge via multi-hop associations, enabling more effective retrieval and integration of KG facts and textual passages. The overall process is illustrated in Figure 1, consisting of three main modules:
1) KG-Guided Spreading Activation. Starting from query-relevant entities, we iteratively expand semantically related triples to build a task-specific subgraph $\mathcal{G}_{q}$, which is then summarized for downstream use.
@@
Details of all prompts used in the framework are in Appendix D.1.
3.3 Preparation: KG Preprocessing
We prepare our KG $\mathcal{G}$ by extracting and preprocessing triples from Wikidata5M [27], a large-scale knowledge graph dataset derived from Wikidata and Wikipedia. Wikidata5M spans multiple domains, making it suitable for open-domain QA tasks. The KG used in KG-Infused RAG is defined as:
where $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{D}_{e}$ denote the set of entities, relations, and entity descriptions, respectively. We preprocess Wikidata5M to meet our requirements for the KG, resulting in Wikidata5M-KG with approximately ~21M triples. The preprocessing details and statistics are in Appendix B.
@@ 3.4 KG-Guided Spreading Activation
We first retrieve the top-$k_{e}$ entities most relevant to the query $q$ as the starting points for spreading activation over the KG. Each entity $e_{i}$ is associated with a textual description $d_{i}\in\mathcal{D}_{e}$. The similarity between the query and each entity is computed via inner product:
@@
We then select the top-$k_{e}$ entities whose descriptions have the highest similarity scores:
$E_{q}^{0}=\{e_{i}\in\mathcal{E}\mid d_{i}\in\text{TopK}_{k_{e}}(\mathcal{D}_{e};\text{sim}(q,d_{i}))\}$ (4)
Here, $\text{TopK}_{k_{e}}(\mathcal{D}_{e};\text{sim})$ denotes the set of $k_{e}$ descriptions in $\mathcal{D}_{e}$ with the highest similarity to $q$, and $e_{i}$ is the entity associated with $d_{i}$. The resulting entity set $E_{q}^{0}=\{e_{1},e_{2},\dots,e_{k_{e}}\}$ initializes the spreading activation process over the KG.
@@
Starting from the seed entities $E_{q}^{0}$, we perform iterative spreading activation over $\mathcal{G}$, where each round $i$ expands from the currently activated entities $E_{q}^{i}$ to retrieve new query-relevant triples. Each round consists of:
@@
    Triple Selection. For each entity in the current set $E_{q}^{i}$, we retrieve its 1-hop neighbors from $\mathcal{G}$. An LLM is prompted to select a subset of triples relevant to the query $q$. The selected triples for round $i$ are denoted as $\{(e,r,e^{\prime})_{i}\}$, where $e$, $r$, and $e^{\prime}$ denote the head entity, relation, and tail entity in a triple respectively.
@@
    Activation Memory Construction and Updating. We maintain an Activation Memory =$\mathcal{M}=\left\{\mathcal{G}_{q}^{act},E_{q}^{act}\right\}$ to track the multi-round activation process, where the subscript $act$ denotes the activation stage. Here, $\mathcal{G}_{q}^{act}$ is the accumulated subgraph of all query-relevant triples retrieved so far, and $E_{q}^{act}$ is the set of entities activated up to the current round. This memory helps aggregate knowledge while preventing redundant reactivation of entities. At each round $i$, given the newly retrieved triples $\mathcal{G}_{q}^{i}$ and the associated entity $E_{q}^{i}$, we update the memory as follows:
@@
    Next-Round Activation Entities. At each round $i$, we determine the next-round entity set $E_{q}^{i+1}$ from the current triples $\mathcal{G}_{q}^{i}=\{(e,r,e^{\prime})_{i}\}$ by extracting their tail entities $\{e^{\prime}\}$ as candidates for further activation. To prevent revisiting entities, we exclude those already in the activated entity set $E_{q}^{act}$. Formally,
@@
The activation process terminates when either the predefined maximum number of activation rounds $k$ is reached or no new entities remain, i.e., $E_{q}^{i+1}=\emptyset$. This iterative expansion progressively retrieves query-relevant triples from $\mathcal{G}$, covering paths from 1-hop to $k$-hop neighborhoods.
@@
We prompt the LLM to summarize the expanded KG subgraph $\mathcal{G}_{q}^{act}$, accumulated in the Activation Memory $\mathcal{M}$, into a natural language summary $\mathcal{S}_{\mathcal{G},{q}}^{act}$. This summarization serves two purposes: (1) It condenses discrete graph-structured facts into a coherent natural language narrative, revealing semantic paths and concept interactions underlying the query. (2) It converts structured knowledge into natural language, making it more accessible and usable for the LLM in subsequent stages.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) improves factual accuracy by grounding responses in external knowledge. However, existing RAG methods either rely solely on text corpora and neglect structural knowledge, or build ad-hoc knowledge graphs (KGs) at high cost and low reliability. To address these issues, we propose KG-Infused RAG, a framework that incorporates pre-existing large-scale KGs into RAG and applies spreading activation to enhance both retrieval and generation. KG-Infused RAG directly performs spreading activation over external KGs to retrieve relevant structured knowledge, which is then used to expand queries and integrated with corpus passages, enabling interpretable and semantically grounded multi-source retrieval. We further improve KG-Infused RAG through preference learning on sampled key stages of the pipeline. Experiments on five QA benchmarks show that KG-Infused RAG consistently outperforms vanilla RAG (by 3.9% to 17.8%). Compared with KG-based approaches such as GraphRAG and LightRAG, our method obtains structured knowledge at lower cost while achieving superior performance. Additionally, integrating KG-Infused RAG with Self-RAG and DeepNote yields further gains, demonstrating its effectiveness and versatility as a plug-and-play enhancement module for corpus-based RAG methods.  [^1] Our code and data are available at https://github.com/thunlp/KG-Infused-RAG
@@ 1 Introduction
Large language models (LLMs) have shown strong performance in question answering tasks (Hurst et al., 2024; Grattafiori et al., 2024; Yang et al., 2024), but remain prone to factual errors and hallucinations (Mallen et al., 2023; Min et al., 2023; Gao et al., 2023b). Retrieval-Augmented Generation (RAG) mitigates these issues by grounding generation in external knowledge sources (Lewis et al., 2020; Trivedi et al., 2023; Asai et al., 2024; Wen et al., 2024). Existing RAG methods can be divided into two main categories: (1) corpus-based approaches that retrieve passages from large-scale text corpora (Gao et al., 2023b; Trivedi et al., 2023; Asai et al., 2024), providing broad knowledge coverage but neglecting important structural information, and (2) KG-based methods (Edge et al., 2024; Guo et al., 2024) that use LLMs to construct ad-hoc knowledge graphs (KGs) from the corpus to enhance the retrieval process, but are limited by retrieval efficiency and the limited amount of information in the constructed KGs. However, both approaches fail to leverage existing large-scale curated KGs, which could enhance both the coverage and efficiency of retrieval.
To bridge these gaps, we propose Knowledge Graph Infused Retrieval-Augmented Generation (KG-Infused RAG), a framework that leverages pre-existing, large-scale KGs to enhance retrieval and generation. At its core, KG-Infused RAG employs spreading activation (Collins & Loftus, 1975), a concept from cognitive psychology in which activation propagates from a central concept to related ones in a semantic network. In RAG, this idea allows LLMs to simulate spreading activation by retrieving structured knowledge from KGs, activating key facts, and progressively accessing query-relevant information. By exploiting KG structure, KG-Infused RAG links related entities and facts often missed by corpus-based methods, leading to more accurate and context-aware retrieval.
Furthermore, KG-Infused RAG eliminates the need for ad-hoc KG construction via LLMs (Edge et al., 2024; Guo et al., 2024; Zhu et al., 2025), which are computationally expensive and prone to factual inaccuracies. Instead, KG-Infused RAG directly utilizes pre-existing KGs, ensuring efficient and reliable retrieval. This use of large-scale KGs enhances retrieval coverage and interpretability, providing a more structured and accurate foundation for downstream generation.
@@
KG-Infused RAG consists of three stages: (1) retrieving relevant KG facts via spreading activation, (2) expanding the query with these facts to improve corpus retrieval, (3) generating answers with passages enriched by KG facts. This multi-source retrieval combines KG facts and corpus passages, leveraging KG structure to guide spreading activation. By aligning the retrieval process with spreading activation, KG-Infused RAG enables more accurate and fact-grounded QA. Compared with KG-based methods such as GraphRAG and LightRAG, KG-Infused RAG avoids costly LLM-based KG construction, achieves more efficient inference, and delivers stronger performance. Additionally, it can be integrated into corpus-based RAG methods like Self-RAG (Asai et al., 2024) to further boost performance.
@@
    We introduce KG-Infused RAG, a framework that incorporates pre-existing KGs into RAG to apply spreading activation, enabling LLMs to retrieve relevant structured knowledge from KGs.
@@
    Experiments on five QA benchmarks show that KG-Infused RAG consistently improves performance, with absolute gains of 3.9% to 17.8% over vanilla RAG.
@@
    KG-Infused RAG outperforms GraphRAG and LightRAG in both efficiency and performance.
@@
    Plug-in application: KG-Infused RAG can be integrated into corpus-based RAG methods (e.g., Self-RAG and DeepNote) as a plug-and-play module.
@@ 2.1 Multi-Turn Retrieval in Retrieval-Augmented Generation
Single-turn retrieval followed by generation often yields insufficient evidence for complex questions, such as multi-hop QA, which requires multiple retrieval steps and the integration and reasoning of retrieved information (Gao et al., 2023b; Trivedi et al., 2023).
@@
Query rewriting reformulates ambiguous or complex queries into clearer, more effective ones (Ma et al., 2023; Mao et al., 2024). Self-Ask (Press et al., 2023) and IRCoT (Trivedi et al., 2023) enhance multi-hop reasoning by decomposing complex queries into sub-questions for iterative retrieval. Query expansion, such as adding keywords or pseudo-documents, enriches the original query to improve retrieval (Wang et al., 2023; Jagerman et al., 2023; Lei et al., 2024). HyDE (Gao et al., 2023a) generates hypothetical documents from the query to improve retrieval relevance. Unlike these methods that rely solely on an LLM’s parametric knowledge, KG-Infused RAG incorporates factual signals from a knowledge graph (KG) to ground query expansion, KG-Infused RAGenables more targeted and evidence-grounded retrieval.
@@
Recent RAG methods extend single-turn retrieval to multi-turn, collecting richer and more targeted evidence (Gao et al., 2023b). IRCoT interleaves retrieval with chain-of-thought reasoning steps, while FLARE (Jiang et al., 2023) and Self-RAG (Asai et al., 2024) dynamically decide when and what to retrieve based on generation confidence or intermediate outputs. These methods enable incremental evidence accumulation. Similarly, KG-Infused RAG gathers additional evidence through multi-turn spading activation and KG-based query expansion, followed by integration and refinement of retrieved information.
@@ 2.2 Knowledge Graphs Augmented Question Answering
Knowledge graphs (KGs) provide structured representations of entities and their relationships, which help reduce hallucinations, improve reasoning interpretability, and enhance retrieval and generation in RAG systems (Baek et al., 2023; Xu et al., 2024; Wen et al., 2024; Guo et al., 2024). MindMap (Wen et al., 2024) and Xu et al. (2024) use KGs as retrieval sources in domains like medical and customer service QA, achieving better retrieval accuracy, answer quality, and transparency. In contrast, GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2024), and KG2RAG (Zhu et al., 2025) generate ad-hoc KGs from the corpus or retrieved passages to organize entity relationships and improve answer coherence for complex questions. However, this requires significant computational resources for KG construction, making it impractical for large-scale corpora. KG-Infused RAG, on the other hand, directly use pre-existing, large-scale KGs, avoiding the costly construction process while providing reliable, structured, and board external knowledge.
@@ 3.1 Task Definition
Let $\mathcal{G}_{q}\subset\mathcal{G}$ denote the retrieved KG subgraph, consisting of query-relevant triples, and $\mathcal{P}_{q}\subset\mathcal{D}_{c}$ represent the corresponding passages that are retrieved. Considering the KG as a semantic network of concept-level associations, the entities within $\mathcal{G}_{q}$ act as seeds, facilitating the propagation of activation throughout the KG. Formally, the task can be framed as:
$\text{answer}\sim\text{LLM}\left(q,\mathcal{G}_{q},\mathcal{P}_{q}\right)$ (1)
@@ 3.2 Framework Overview
Figure 1: Overview of KG-Infused RAG. The framework consists of three modules: (1) KG-Guided Spreading Activation, (2) KG-Based Query Expansion, and (3) KG-Augmented Answer Generation, which together enable interpretable, fact-grounded multi-source retrieval.
We propose KG-Infused RAG, a framework inspired by the cognitive process of spreading activation, where activation spreads from a central concept to semantically related ones. By leveraging the rich semantic structure of pre-existing, large-scale KGs, KG-Infused RAG extends relevant knowledge via multi-hop associations, enhancing the retrieval and integration of both KG facts and textual passages, without the need for ad-hoc KG construction. The overall process is illustrated in Figure 1, consisting of three main modules:
1) KG-Guided Spreading Activation. Starting from query-relevant entities, we iteratively expand related triples to form a task-specific subgraph $\mathcal{G}_{q}$, which is summarized for downstream use.
@@
We prepare our KG $\mathcal{G}$ by extracting and preprocessing triples from a large-scale knowledge graph dataset, Wikidata5M. Details of the preprocessing are provided in Appendix A. Details of all prompts used in the framework are in Appendix C.1.
@@ 3.3 KG-Guided Spreading Activation
To initialize the spreading activation process, we first retrieve the top-$k_{e}$ entities most relevant to the query $q$. Each entity $e_{i}$ is associated with a textual description $d_{i}\in\mathcal{D}_{e}$, and we compute the similarity between the query and each description using inner product:
@@
The top-$k_{e}$ entities with the highest similarity scores are selected:
$E_{q}^{0}=\{e_{i}\mid d_{i}\in\text{TopK}_{k_{e}}(\mathcal{D}_{e};\text{sim}(q,d_{i}))\}$ (3)
where $\text{TopK}_{k_{e}}(\mathcal{D}_{e};\text{sim})$ denotes the set of $k_{e}$ descriptions in $\mathcal{D}_{e}$ with the highest similarity to $q$. The resulting entity set $E_{q}^{0}=\{e_{1},e_{2},\dots,e_{k_{e}}\}$ initializes the spreading activation process over the KG.
@@
The spreading activation process begins with the seed entities $E_{q}^{0}$ and proceeds iteratively over $\mathcal{G}$. In each round $i$, activation expands from the currently activated entities $E_{q}^{i}$, retrieving additional query-relevant triples. Each round consists of the following steps:
@@
    Triple Selection. At each round, for every entity in the current set $E_{q}^{i}$, we retrieve its 1-hop neighbors from $\mathcal{G}$. An LLM is then prompted to select a subset of triples relevant to the query $q$. The selected triples for round $i$ are denoted as $\{(e,r,e^{\prime})_{i}\}$, where $e$, $r$, and $e^{\prime}$ represent the head entity, relation, and tail entity in a triple, respectively.
@@
    Activation Memory Construction and Updating. We maintain an Activation Memory $\mathcal{M}=\left\{\mathcal{G}_{q}^{act},E_{q}^{act}\right\}$, which tracks the multi-round activation process. The subscript actact denotes the activation stage. Here, $\mathcal{G}_{q}^{act}$ is the accumulated subgraph of all query-relevant triples retrieved up to the current round, and $E_{q}^{act}$ is the set of entities activated thus far. This memory serves to aggregate knowledge and prevent redundant reactivation of entities. At each round $i$, given the newly retrieved triples $\mathcal{G}_{q}^{i}$ and the associated entities $E_{q}^{i}$, we update the memory as follows:
@@
    Next-Round Activation Entities. At each round $i$, we determine the next-round entity set $E_{q}^{i+1}$ from the current triples $\mathcal{G}_{q}^{i}=\{(e,r,e^{\prime})_{i}\}$ by extracting their tail entities $\{e^{\prime}\}$ as candidates for further activation. To prevent revisiting entities, we exclude those already present in the activated entity set $E_{q}^{act}$. Formally,
@@
The activation process terminates when either the predefined maximum number of activation rounds $k$ is reached or no new entities remain, i.e., when $E_{q}^{i+1}=\emptyset$. This iterative process progressively retrieves query-relevant triples from $\mathcal{G}$, spanning paths from 1-hop to $k$-hop neighborhoods.
@@
We prompt the LLM to summarize the expanded KG subgraph $\mathcal{G}_{q}^{act}$, accumulated in the Activation Memory $\mathcal{M}$, into a natural language summary $\mathcal{S}_{\mathcal{G},{q}}^{act}$. This summarization serves two purposes: (1) It condenses discrete graph-structured facts into a coherent natural language narrative, uncovering semantic paths and concept interactions underlying the query. (2) It converts structured knowledge into natural language, making it more accessible and usable for the LLM in the subsequent stages.